OpenStudy (anonymous):

Solve the given equation and indicate the number of values of θ sach that 0≤θ<2π ,that satisfy the iquation; 2tanθsinθ-tanθ=0

OpenStudy (anonymous):

\[θ = 0\] \[θ = \pi/6\] \[θ = (5 \pi)/6\] \[θ = \pi\]